OpenStudy (faithspring):

Solve this expression (x^3) ^ -5

rebeccaxhawaii (rebeccaxhawaii):

1/x^15

rebeccaxhawaii (rebeccaxhawaii):

3 x 5 = 15

OpenStudy (faiqraees):

when a power is being powered to a certain degree the powers are multiplied \[(a^m)^n =a^{m*n}\]

OpenStudy (gabbyalicorn):

(x^3)^−5 =(x^3)^−5 =1/x^3*x^3*x^3*x^3*x^3 =1x^15